SIAS Web Interface
SIAS Date: 2024-05-18Time of Printout: 19:13
May 2024 (HST)
SunMonTueWedThuFriSat
 
18
 
To ->.
Go to
[hide]
STATUSK1K2
Telescopes
Keck 1 READY
Keck 2 READY
Scheduled
Instruments
OSIRIS-LGS
(osiris3)
READY
PCS
(???)
READY
NIRC2-LGS
(nirc6)
READY
Standby
Instruments
MOSFIRE
READY
KPF
READY
HIRESr
READY
NIRES
READY
Last Update Time: 2024-05-18 19:13 (HST)
SUPPORTK1K2
Observers[none] [Bentley, Chu, Do, Ghez, Hosek, Do, Chu][none] [Do, Granados, Paugnat]
SAJim L.Carlos A.
OAJohn P.
summit
Rita M.
rem-ops
SWOCMatthew B.
EEOCBrett S.
Summit Support Lead Zach Tadio Christopher Funada
Who System Description Last Changed Job Done Comments
Initials Time
SupportLRIS**HIGH PRIORITY** Fill LN after arriving in the morning.
2024-05-18 11:01
DB 11:00complete 
SupportKCWI**HIGH PRIORITY** Fill RED Dewar LN *AFTER* 9:30am *CONFIRM SUCCESSFULL FILL COMPLETED AT KCWI INSTRUMENT PLC TOUCHSCREEN* "Red fill complete, LN2 detected in vent"
2024-05-18 11:42
CF 11:45Red fill complete 
Support
2024-02-02 12:34
 00:00 
SupportESI**HIGH PRIORITY** Fill LN on arrival -- Cancelled
2024-05-18 11:42
CF 11:45jlyke--ESI off-sky until 2024-05-30 -- Cancelled 
SupportNIRES**HIGH PRIORITY** Fill Main Tank LN on arrival
2024-05-18 13:24
CF 12:30Fill completed 
SupportK1 FacilityNaz deck Tool Board and Binder checkout
2024-05-18 09:17
ZT 09:15complete 
SupportK2 FacilityNaz deck Tool Board and Binder checkout
2024-05-18 11:42
CF 11:45Morning check complete 
SupportK2 FacilityFacility reset to LOCAL control.
2024-05-18 11:43
CF 11:45Complete 
SupportK1 FacilityFacility reset to LOCAL control.
2024-05-18 12:23
ZT 12:15complete 
SupportKPFBy 10:00 am verify the 180l dewar in the basement is at least 1/3 full
2024-05-18 15:47
DB 15:45dewar changed out 
SupportKPFAfter 12:00pm confirm both Red and Green KPF dewars are at least 145 lbs on the K2 control touchscreen. Reset any timeouts
2024-05-18 12:48
DB 12:45Green weight = 150 Red Weight = 150 
SupportDome TempsAfter 10:00 AM, Set both Dome temps to the MKWS forecasted 8:00 PM temperatures, found here: http://hokukea.soest.hawaii.edu/forecast/mko/ Please log temps in comments.
2024-05-18 16:04
CF 12:00Dome temps set to 31.1° (No changes) 
SupportDEIMOSFill LN in afternoon.*CONFIRM SUCCESSFULL FILL COMPLETED AT DEOMOS INSTRUMENT PLC TOUCHSCREEN*
2024-05-18 15:50
CF 15:45Afternoon fill complete Liquid closed Hose disconnected Bandbreak on 
SupportK2Reconfigure K2:
  • change OBaffles from Retracted to Deployed;
  • 2024-05-18 11:44
    CF 11:45Obaffles deployed DEIMOS weights removed Horizon: +0.8up, -0.5dn Zenith: +1.0up, -0.5dn 
    SupportPCS2Perform checkout on the PCS System
    2024-05-18 16:03
    CF 16:00Checkout complete 
    SupportK1 FacilityNaz deck Tool Board and Binder checkout
    2024-05-18 12:37
    ZT 12:30complete 
    SupportK2 FacilityNaz deck Tool Board and Binder checkout
    2024-05-18 16:14
    CF 16:15Afternoon check complete 
    SupportK1LGSK1 LGS Afternoon Checkout
    2024-05-18 15:47
    ZT 15:45k1 laser check out complete, laser left in ready state, all is well 
    Laser TechK2LGSK2 Laser Checkout
    2024-05-18 16:02
    CF 15:45Laser checkout is complete Laser propagated at 30,60,90 Laser left in READY state 
    SupportOSIRIS-LGSK1 Instrument & Guider READY
    2024-05-18 15:49
    ZT 12:30OSIRIS-LGS Instrument guider check and test all complete 
    SupportNIRC2-LGSK2 Instrument & Guider READY
    2024-05-18 16:14
    CF 16:15Guider check and test all complete 
    SupportPCSK2 Instrument & Guider READY
    2024-05-18 16:15
    CF 16:15Guider check complete 
    SupportMOSFIREK1 Standby Instrument READY
    2024-05-18 15:48
    ZT 12:30MOSFIRE Instrument guider check and test all complete 
    SupportKPFK1 Standby Instrument READY
    2024-05-18 15:48
    ZT 12:30KPF Instrument guider check and test all complete 
    SupportHIRESrK1 Standby Instrument READY
    2024-05-18 15:48
    ZT 12:30HIRES Instrument guider check and test all complete 
    SupportNIRESK2 Standby Instrument READY
    2024-05-18 16:15
    CF 16:15Guider check and test all complete 
    SupportK1 TelescopeK1 Telescope READY
    2024-05-18 16:10
    ZT 16:15k1 telescope parked dark and released in dome flats position 
    SupportK2 TelescopeK2 Telescope READY
    2024-05-18 16:15
    CF 16:15K2 parked dark in dome flats position 
    SupportKPFAt around 2:30 pm, make sure the dewar in the basement has enough for tomorrows fill or at least 1/3 full
    2024-05-18 15:48
    DB 12:45new dewar installed on KPF 
    SupportLRISFill LN after telescope initialization
    2024-05-18 16:09
    ZT 16:15complete 
    SupportESIFill LN after telescope initialization -- Cancelled
    2024-05-18 16:03
    CF 16:00jlyke--ESI off-sky until 2024-05-30 -- Cancelled 
    Last completed
    TBAD task:
    K1: 2024-05-16
    (2 days ago)
    K2: 2024-05-16
    (2 days ago)

    The server is: mysqlserver [readonly mode]

    The datebook is: datebook

    SIAS version 5.3.1